# Effort-aware Fairness: Incorporating a Philosophy-informed, Human-centered Notion of Effort into Algorithmic Fairness Metrics

### v1

@@ Introduction
Next, we must decide on a more convenient fairness metric (e.g., group or individual fairness) to elicit fairness judgements from laypeople and measure the impact of effort-related features on such judgements. We choose individual fairness over group fairness because individual fairness evaluation minimizes the effect of participants’ cognitive load and mathematical ability as potential confounders. If we validated the role of effort in people’s group fairness perception instead, we could not simply ask people to make fair decisions for two individuals, but for many individuals simultaneously from varying demographics to compute group parity metrics, increasing their cognitive load and making their fairness judgement more sensitive to non-substantive factors like visualization techniques (Van Berkel et al. 2021). A group fairness setting also assumes that people can compute group statistics (percentages) in their heads if they want to make group-wise fair decisions, which is not true even for college-level laypeople (Jacobs Danan and Gelman 2018).
@@
Individual fairness is formulated as “similar individuals are treated similarly” (Dwork et al., 2012). To evaluate whether a decision-making process (with or without AI) satisfies the individual fairness criterion, the probabilistic formulation of individual fairness might be simplified into two simple, more deterministic steps: (1) (Overall) Input-space distance: evaluate pairwise distance (based on features in the input space) to find pairs of highly “similar” individuals; (2) Output-space distance: in each pair of two highly similar individuals, check if their two outcomes are similar enough.
@@
Our main question is whether people consider pairwise distance in terms of a feature’s aggregate value (“aggregate distance”) more or in terms of its trajectory (“trajectory distance”) more during each of the two aforementioned steps in the individual fairness evaluation pipeline. We design two types of trajectory: increasing vs. decreasing, which means positive vs. negative velocity in a non-cumulative formulation (e.g., annual income) as shown in Figure 1, and is mathematically equivalent to positive vs. negative acceleration in a cumulative formulation (e.g., cumulative income since the first year). We show participants a non-cumulative (annual) feature view because many laypeople would not be able to distinguish positive vs. negative second-order derivative (convex vs. concave curvature) in a cumulative view. Therefore, “trajectory” in this experiment corresponds to the cumulative-view “acceleration” term, a cornerstone in our philosophy-informed formulation of effort. We draw real-world (anonymized) samples of defendants/households from CLUE (with past arrests as an effort feature) and SHED (with income as an effort feature). Our research questions:
@@
The two main dependent variables are overall input-space distance and output-space distance. Overall input-space distance is measured by asking participants, after they have rated the trajectory distance and aggregate distance for a pair, to rate, on the same Likert scale, how different the two decision subjects are overall. Regarding output-space distance, in each set of $5$ decision subjects, $1$ of them is the reference subject ($K$) and the other $4$ are comparison subjects $(A,B,C,D)$. We trained an AI model to predict the risk score (violent recidivism risk in CLUE; balance payment failure risk in SHED) for the reference subject $K$, which each participant can either accept or correct with their own perceived risk score for $K$. Then, we ask the participant to provide a“fair” risk score for each of the $4$ comparison subjects $(A,B,C,D)$, considering the risk score they fixed for $K$ and the input-space features (arrests in CLUE, income in SHED) of the 4 pairs $(K,A)$, $(K,B)$, $(K,C)$, and $(K,D)$, which by design have (similar trends, similar totals), (different trends, similar totals), (similar trends, different totals), and (different trends, different totals), respectively.
Figure 1: Example SHED set of 1 reference subject (K) and 4 subjects to be compared to K. A: similar trajectory/aggregate; B: different trajectory, similar aggregate; C: similar trajectory, different aggregate; D: different trajectory/aggregate. All households’ incomes are real SHED records, except that we slightly perturb the first time step of K to get A.
@@
where $\tilde{M}$ presents participants’ mental predictive model of the risk scores. To test if our findings will be robust across different AI application contexts and different effort-related features, each participant will be randomly assigned to one of the two between-subjects conditions: CLUE (where the effort-related, input-space feature is defendants’ number of previous arrests, and the output-space feature is defendants’ risk of committing a violent crime if released on bail) vs. SHED (where the effort-related, input-space feature is households’ income, and the output-space feature is households’ risk of not pay back a monthly credit balance).
@@
We show interfaces for the main types of questions in the experiment in Figure 2 (SHED condition) and Figure 4 (CLUE condition) in Appendix E. Since the answers to these questions are highly subjective, to incentivize participants to think carefully and give informedly truthful opinions, we apply the Bayesian Truth Serum (BTS) method introduced by Prelec (2004) (details in Appendix E). The BTS method was validated and applied in fields ranging from Psychology (John, Loewenstein, and Prelec 2012) and Marketing (Weaver and Prelec 2013) to AI (Witkowski and Parkes 2012) and HCI (Miller, Bailey, and Kirlik 2014). Therefore, we use a simplified version of BTS on our most subjective type of MCQ: overall input-space distance, with a \$$\$5$ “Honesty bonus” for the top $10\%$ participants in BTS score.
For the other type of dependent variable question (output-space distance), since this question asks people to estimate a numeric “fair” risk score, BTS (designed for MCQ) is not suited. Our alternative incentive mechanism is a \$$\$5$ “Rationale bonus” for the top $10\%$ participants who write the most detailed and persuasive rationales to justify their risk scores. As people write down justifications, they have more time to think and potentially re-calibrate their risk scores.
We apply the following filters when recruiting Prolific participants: approval rate of at least $99\%$, Bachelor’s degree, U.S. residency, and English fluency. Considering the median study completion time of approx. $42$ minutes, we pay each participant \$8.45, corresponding to a median pay rate of \$12 per hour. After excluding $6$ participants who failed $2$ out of $3$ attention checks pursuant to Prolific’s policy, our final data include $149$ samples.999144 regular participants and $5$ timed-out participants, who exceeded the maximum time allowed by Prolific but still finished the survey on Qualtrics
This experiment has been approved by our Institutional Review Board (IRB). We also pre-registered our experimental hypotheses and analyses on Jan $10$, $2025$ before recruiting Prolific subjects on Jan $11$, $2025$ to collect data. [^10] Anonymized proof of pre-registration: https://aspredicted.org/4qg6-zpfm.pdf
Figure 2: Example human subjects experiment interfaces to measure: 1/ trajectory distance (upper left); 2/ aggregate distance (upper right); 3/ overall input-space distance (lower left); 4/ output-space distance (lower right). Condition: SHED.
@@
To address RQ1, we compute Spearman correlation coefficients to examine the relationships between trajectory distance and overall input-space distance, as well as between aggregate distance and overall input-space distance. Spearman correlation is appropriate here because it does not assume linearity and is robust to non-normal distributions and outliers, which are suitable for ordinal data. We then perform one-sample t-tests to determine if these correlations are significantly different from 00. If both correlations are significant, we use Hotelling’s t-test (May and Hittner 1997) and Steiger’s Z-test (Meng, Rosenthal, and Rubin 1992) to compare the strengths of these correlations. Hotelling’s t-test is well-suited for assessing differences in dependent correlations, as it accounts for the shared dependent variable. However, it is known to be overly powerful, which can come at the expense of a higher Type I error rate in some scenarios (May and Hittner 1997). Steiger’s Z-test, on the other hand, provides a more generalizable framework and is less sensitive to deviations from normality by implementing a Fisher’s z transformation of sample correlation coefficients. This makes Steiger’s Z-test particularly valuable for larger datasets where the normality assumption may not hold perfectly. By using both tests, we can balance the strengths and limitations of each approach, providing more reliable comparisons of correlated correlation coefficients. To address concerns related to multiple testing and reduce the likelihood of Type I errors, we apply the Bonferroni correction to adjust the p-values. To further validate these findings, we employ a bootstrapping procedure to estimate the confidence intervals of the differences between the two correlation coefficients. Moreover, we estimate a linear regression model:
@@
For RQ3, we perform a regression analysis to test if the between-subjects variable moderate the relationships examined in RQ1 and RQ2. We estimate the following models:
@@
In both the CLUE and SHED conditions, trajectory distance exhibits stronger correlations with overall input-space and output-space distances compared to aggregate distance, as shown in Table 2 (Appendix F). Under the CLUE condition, trajectory distance shows a strong correlation with overall input distance ($r_{\text{traj},\text{input}}=0.7056$) and a moderate correlation with output distance ($r_{\text{traj},\text{output}}=0.4218$), whereas aggregate distance has weaker correlations ($r_{\text{aggr},\text{input}}=0.5022$ for input and $r_{\text{aggr},\text{output}}=0.2089$ for output). A similar pattern is observed in the SHED condition, where trajectory distance strongly correlates with overall input distance ($r_{\text{traj},\text{input}}$ = 0.7881) and moderately with output distance ($r_{\text{traj},\text{output}}$ = 0.4986), while aggregate distance shows weaker correlations ($r_{\text{aggr},\text{input}}=0.6547$ for input and $r_{\text{aggr},\text{output}}=0.2507$ for output).
[…]
@@
To further validate our findings, we implement a bootstrapping procedure in Table 4 to estimate the confidence interval of the difference between the two correlation coefficients. The $95\%$ confidence interval of the differences between $r_{\text{traj},\text{input}}$ and $r_{\text{aggr},\text{input}}$ and is $[-0.2780,-0.1280]$ for CLUE and $[-0.1950,-0.0730]$ for SHED, with neither interval containing 00. We can conclude that trajectory distance correlates more strongly with overall input-space distance than aggregate distance does. The $95\%$ confidence interval of the differences between $r_{\text{traj},\text{output}}$ and $r_{\text{aggr},\text{output}}$ is $[-0.3000,-0.1280]$ for CLUE and $[-0.3330,-0.1610]$ for SHED. Both intervals also exclude 00, confirming that trajectory distance correlates more strongly with overall output-space distance than aggregate distance.
[…]
@@
After each question on overall input-space distance, we have an optional rationale question where participants may select all factor(s) that influence the overall input-space distance they assigned to a pair of defendants/households, including trajectory/trend, aggregate/total, and “Other (please specify)”. Participants select trajectory more often than aggregate in both CLUE (80.6% $>$ 71.4%) and SHED (75.2% $>$ 60.0%) conditions, which further affirms our quantitative results. We perform thematic coding inductively on the 55 free-text responses (rationales) that follow the third choice, “Other (please specify below)” (1.9% in CLUE, 4.6% in SHED), and come up with 9 themes (after filtering out 4 vague rationales). Although people are expected to specify “Other” factors, the majority of rationales still is still about trajectory and/or aggregate, as shown in the following three themes (note that we concatenate the Response ID after each excerpt): 1/ Both trajectory and aggregate (23 rationales, e.g., “These cases were very different in terms of trajectory, only meeting once. However, their aggregate was the same, so I chose an option sort of in between these, but closer to the similar side for my overall thoughts.”); 2/ Only trajectory (12 rationales, e.g., “The total arrests aren’t similar but the way they are both trending upwards makes them similar.”); 3/ Only aggregate (3 rationales, e.g., “I ultimately landed with dissimilar via the aggregate of income”).
New themes that influence how people evaluate overall input-space distance include: 4/ Concern about unobserved factors (4 rationales, e.g., “Possible retirement or loss of job by B”); 5/ First time step (3 rationales, e.g., “Similar/near equal high amount of arrest to start with.”); 6/ Last time step (2 rationales, e.g., “In addition to the vastly different trajectories, it is important to consider the endpoint of the graph, as it is at this time that the decision will be made, such as by a bank.”); 7/ Magnitude of acceleration (2 rationales, e.g., “Jumps in increased or decreased crimes.”). Two other rationales are unique enough to each define their own theme: 8/ Comparing to an average person: “Although the total arrest times have a significant different, the number of total arrests for both people is quite a bit and more than an average person”; 9/ Intersection: “The intersection of their trajectories”.
@@
One inertia metric that is relevant to both criminal justice and personal finance is childhood poverty, as explained earlier. However, for adults in our two real-world datasets, information on adults’ childhood socio-economic background is not available, especially not at the individual level. Therefore, to quantify inertia empirically, we may use race as an individual-level but imperfect proxy for childhood poverty (or inertia) in the U.S. as the National Center for Education Statistics (NCES) found that childhood poverty rate differs substantially across races (e.g., White: $13\%$, Asian: $14\%$, Pacific Islander: $25\%$, American Indian: $36\%$, Black: $39\%$) in 2012.111111https://nces.ed.gov/programs/coe/indicator/cce/family-characteristics. Figure 4. Accessed Jan 19, 2025 To obtain $m$ per race, we scale these poverty rates by the maximum rate (39%) so that $m$ spreads across $[0,1]$.
Given four equidistant time steps $\{0,1,2,3\}$ (assuming $\Delta t=1$), we model cumulative arrests in CLUE (or cumulative income, in multiples of \$$\$10k$ in SHED) at each time step as $X_{0}$, $X_{1}$, $X_{2}$, $X_{3}$. Their velocities are $V_{0}$, $V_{1}$, $V_{2}$, where $V_{i}=\frac{X_{i+1}-X_{i}}{\Delta t}=X_{i+1}-X_{i}$. Their accelerations are $A_{0}$ and $A_{1}$, where $A_{i}=\frac{V_{i+1}-V_{i}}{\Delta t}=V_{i+1}-V_{i}$. We average these two acceleration terms to get $A_{\text{avg}}=\frac{1}{2}(A_{0}+A_{1})=\frac{1}{2}(v_{1}+v_{2})$. Noting that the effort-related feature is desirable in SHED (income) but undesirable in CLUE (arrests), we compute the individual-level (good) effort, $E$, with context adaptations:
@@
The function $\sigma(z)=\frac{1}{1+e^{-z}}$ monotonically maps real values from (,)($\infty$,$\infty$)( - ∞ , ∞ ) to the $(0,1)$ range, serving two purposes. First, it allows $\sigma(A_{\text{avg}})$ to only obtain one possible sign. More specifically, considering the desirable effort feature, income, in SHED. Note that $m$ is always positive. When the average acceleration term is positive, $E_{\text{good}}=m\cdot A_{\text{avg}}$ will assign better (more positive) effort to those with higher inertia (from more disadvantaged backgrounds). However, when the average acceleration term is negative, $E_{\text{good}}=m\cdot A_{\text{avg}}$ will assign worse (more negative) effort to those with higher inertia, which contradicts our intention (given the same acceleration, more disadvantaged people should be credited for better efforts). The sigmoid function eliminates the possibility of a negative sign and solves this problem. Second, the sigmoid function gathers extreme values closer together so that we have more data points to compute group parity even for subjects with extreme accelerations.
@@
To ensure the aggregate income feature stays in [0,1] like the effort feature and reflects the intuition that the same income gap at lower income (e.g., \$$\$10k$ vs. \$$\$20k$) is more significant than at higher income (e.g., \$$\$110k$ vs. \$$\$120k$), we define an aggregate (sum) function $S(\mathbf{x})$ by applying a modified (scaled and translated) right-hand side of the sigmoid function $\sigma$ on the whole four-year (non-negative) income $X_{3}=\sum_{t=0}^{3}x_{t}$, scaled by $\lambda=\$200k$ so that this aggregate feature is empirically spread out across [0,1] range:
@@
Table 1 shows that according to our EaIF metric, XGBoost significantly less fair than the other two models, and random forest is slightly more fair than logistic regression. This result is consistent across both set of weights (between the effort feature and the most recent cumulative feature).
@@
Binns (2020) found that group fairness and individual fairness are rooted in the same philosophical basis (both are linked to philosophical concepts such as “consistency” and “egalitarianism”). Technical works developed algorithms that satisfy both group fairness and individual fairness, demonstrating the two metrics’ compatibility (Zemel et al. 2013). Therefore, we will also demonstrate how to compute Effort-aware Group Fairness (EaGF) with our same force-based notion of “effort”, which was validated by our human subjects experiment in the individual fairness context.
Given an individual-level effort feature, our goal is to compute effort-aware group parity and observe this metric as a function of numeric effort. Our main steps include 1/ partition decision subjects into similar-effort bins; 2/ compute a demographic group parity metric within each bin; 3/ plot this parity metric against the effort bins. We used demographic features of each household representative who completed the SHED survey to compute group fairness metrics.

### v3

@@ Introduction
Next, we must decide on a more convenient fairness metric (e.g., group or individual fairness) to elicit fairness judgements from laypeople and measure the impact of Effort-related features on such judgements. We choose individual fairness over group fairness because individual fairness evaluation minimizes the effect of participants’ cognitive load and mathematical ability as potential confounders. [^10] If we validated the role of Effort in people’s group fairness perception instead, we could not simply ask people to make fair decisions for two individuals, but for many individuals simultaneously from varying demographics to compute group parity metrics, increasing their cognitive load and making their fairness judgement more sensitive to non-substantive factors like visualization techniques (Van Berkel et al. 2021). A group fairness setting also assumes that people can compute group statistics (percentages) in their heads if they want to make group-wise fair decisions, which is not true even for college-level laypeople (Jacobs Danan and Gelman 2018).
@@
Individual fairness is formulated as ”similar individuals are treated similarly” (Dwork et al., 2012). To evaluate whether a decision-making process (with or without AI) satisfies the individual fairness criterion, the probabilistic formulation of individual fairness might be simplified into two simple, more deterministic steps: (1) (Overall) Input-space distance: evaluate pairwise distance (based on features in the input space) to find pairs of highly ”similar” individuals; (2) Output-space distance: in each pair of two highly similar individuals, check if their two outcomes are similar enough.
@@
Our main question is whether people consider pairwise distance in terms of a feature’s aggregate value (”aggregate distance”) more or in terms of its trajectory (”trajectory distance”) more during each of the two aforementioned steps in the individual fairness evaluation pipeline. We design two types of trajectory: increasing vs. decreasing, which means positive vs. negative velocity in a non-cumulative formulation (e.g., annual income) as shown in Figure 1, and is mathematically equivalent to positive vs. negative acceleration in a cumulative formulation (e.g., cumulative income since the first year). We show participants a non-cumulative (annual) feature view because many laypeople would not be able to distinguish positive vs. negative second-order derivative (convex vs. concave curvature) in a cumulative view. Therefore, ”trajectory” in this experiment corresponds to the cumulative-view ”acceleration” term, a cornerstone in our philosophy-informed formulation of Effort. We draw real-world (anonymized) samples of defendants/households from CLUE (with past arrests as an Effort feature) and SHED (with income as an Effort feature). Our research questions:
@@
The two main dependent variables are overall input-space distance and output-space distance. Overall input-space distance is measured by asking participants, after they have rated the trajectory distance and aggregate distance for a pair, to rate, on the same Likert scale, how different the two decision subjects are overall. Regarding output-space distance, in each set of $5$ decision subjects, $1$ of them is the reference subject ($K$) and the other $4$ are comparison subjects ($A,B,C,D$).
Figure 1: Example SHED set of 1 reference subject (K) and four subjects to be compared to K. A: similar trajectory/aggregate; B: different trajectory, similar aggregate; C: similar trajectory, different aggregate; D: different trajectory/aggregate. All households’ incomes are real SHED records, except that we slightly perturb the first time step of K to get A.
@@
where $\tilde{M}$ presents participants’ mental predictive model of the risk scores.
To test if our findings will be robust across different AI application contexts and different Effort-related features, each participant will be randomly assigned to one of the two between-subjects conditions: CLUE (where the Effort-related, input-space feature is defendants’ number of previous arrests, and the output-space feature is defendants’ risk of committing a violent crime if released on bail) vs. SHED (where the Effort-related, input-space feature is households’ income, and the output-space feature is households’ risk of not paying back a monthly credit balance).
@@
We show interfaces for the main types of questions in the experiment in Figure 2 (SHED condition) and Figure 4 (CLUE condition, Appendix E). Since the answers to these questions are highly subjective, to incentivize participants to think carefully and give informedly truthful opinions, we apply the Bayesian Truth Serum (BTS) method introduced by Prelec (2004) (details in Appendix E). The BTS method was validated and applied in fields ranging from Psychology (John, Loewenstein, and Prelec 2012) and Marketing (Weaver and Prelec 2013) to AI (Witkowski and Parkes 2012) and HCI (Miller, Bailey, and Kirlik 2014). Therefore, we use a simplified version of BTS on our most subjective type of MCQ: overall input-space distance, with a \$5\\$5 ”Honesty bonus” for the top $10\%$ participants in BTS score.
For the other type of dependent variable question (output-space distance), since this question asks people to estimate a numeric ”fair” risk score, BTS (designed for MCQ) is not suited. Our alternative incentive mechanism is a \$5\\$5 ”Rationale bonus” for the top $10\%$ participants who write the most detailed and persuasive rationales to justify their risk scores. As people write down justifications, they have more time to think and potentially recalibrate their risk scores.
We apply the following filters when recruiting Prolific participants: approval rate of at least $99\%$, Bachelor’s degree, U.S. residency, and English fluency. Considering the median study completion time of approximately. $42$ minutes, we pay each participant \$8.45\\$8.45, corresponding to a median pay rate of \$12\\$12 per hour. After excluding $6$ participants who failed $2$ out of $3$ attention checks under Prolific’s policy, our final data include $149$ samples.111111144 regular participants and $5$ timed-out participants, who exceeded the maximum time allowed by Prolific but still finished the survey on Qualtrics
This experiment has been approved by our Institutional Review Board (IRB). We also pre-registered our experimental hypotheses and analyses on Jan $10$, $2025$ before recruiting Prolific subjects on Jan $11$, $2025$ to collect data. [^12] Pre-registration proof: https://aspredicted.org/4qg6-zpfm.pdf
Figure 2: Example human subjects experiment interfaces to measure: (1) trajectory distance (upper left); (2) aggregate distance (upper right); (3) overall input-space distance (lower left); (4) output-space distance (lower right).
@@
To address RQ1, we compute Spearman correlation coefficients to examine the relationships between trajectory distance and overall input-space distance, as well as between aggregate distance and overall input-space distance. Spearman correlation is appropriate here because it does not assume linearity and is robust to non-normal distributions and outliers, which are suitable for ordinal data. We then perform one-sample t-tests to determine if these correlations are significantly different from 0. If both correlations are significant, we use Hotelling’s t-test (May and Hittner 1997) and Steiger’s Z-test (Meng, Rosenthal, and Rubin 1992) to compare the strengths of these correlations. Hotelling’s t-test is well-suited for assessing differences in dependent correlations, as it accounts for the shared dependent variable. However, it is known to be overly powerful, which can come at the expense of a higher Type I error rate in some scenarios (May and Hittner 1997). Steiger’s Z-test, on the other hand, provides a more generalizable framework and is less sensitive to deviations from normality by implementing Fisher’s z transformation of sample correlation coefficients. This makes Steiger’s Z-test particularly valuable for larger datasets where the normality assumption may not hold perfectly. By using both tests, we can balance the strengths and limitations of each approach, providing more reliable comparisons of correlation coefficients. To address concerns related to multiple testing and reduce the likelihood of Type I errors, we apply the Bonferroni correction to adjust the p-values. To further validate these findings, we employ a bootstrapping procedure to estimate the confidence intervals of the differences between the two correlation coefficients. Moreover, we estimate a linear regression model:
@@
For RQ3, we perform a regression analysis to test if the between-subjects variable moderates the relationships examined in RQ1 and RQ2. We estimate the following models:
@@
Our responses to RQ1 and RQ2 are affirmative. In both the CLUE and SHED conditions, trajectory distance exhibits stronger correlations with overall input-space and output-space distances compared to aggregate distance, as shown in Table 2 (Appendix F). Under the CLUE condition, trajectory distance shows a strong correlation with overall input distance ($r_{\text{traj},\text{input}}=0.7056$) and a moderate correlation with output distance ($r_{\text{traj},\text{output}}=0.4218$). In contrast, aggregate distance has weaker correlations ($r_{\text{aggr},\text{input}}=0.5022$ for input and $r_{\text{aggr},\text{output}}=0.2089$ for output). In the SHED condition, trajectory distance strongly correlates with overall input distance ($r_{\text{traj},\text{input}}$ = 0.7881) and moderately with output distance ($r_{\text{traj},\text{output}}$ = 0.4986), while aggregate distance shows mixed correlations with a moderate correlation with input distance ($r_{\text{aggr},\text{input}}=0.6547$) and a weak output correlation ($r_{\text{aggr},\text{output}}=0.2507$).
@@
Table 3 (Appendix F) shows Hotelling’s t-test and Steiger’s Z-test to compare these correlations’ strengths. Across conditions, trajectory distance consistently shows significantly stronger correlations with overall input and output distances than aggregate distance. [^13] In the CLUE condition, the differences in correlations are highly significant, with Hotelling’s t-test yielding statistics of $7.9758$ for input-space and $5.5161$ for output-space distances ($\text{p-value}<0.001$). Steiger’s Z-test further supports these results with statistics of $6.9077$ for input and $5.3461$ for output ($\text{p-value}<0.001$). In the SHED condition, the differences are statistically robust, with Hotelling’s t-test statistics of $7.3360$ (input) and $7.9612$ (output), and Steiger’s Z-test statistics of $6.5541$ (input) and $7.6466$ (output), all significant at $\text{p-value}<0.001$ after Bonferroni correction. To further validate our findings, we implement a bootstrapping procedure in Table 4 (Appendix F) to estimate the confidence interval of the difference between the two correlation coefficients.  [^14] The $95\%$ confidence interval of the differences between $r_{\text{traj},\text{input}}$ and $r_{\text{aggr},\text{input}}$ and is [0.2780,0.1280][-0.2780,-0.1280] for CLUE and [0.1950,0.0730][-0.1950,-0.0730] for SHED, with neither interval containing 0. The $95\%$ confidence interval of the differences between $r_{\text{traj},\text{output}}$ and $r_{\text{aggr},\text{output}}$ is [0.3000,0.1280][-0.3000,-0.1280] for CLUE and [0.3330,0.1610][-0.3330,-0.1610] for SHED. Both intervals also exclude 0.
As shown in Table 5 (Appendix F), after controlling for aggregate distance, trajectory distance remains significantly correlated with overall input-space distance [^15] In CLUE, our regression model estimates $\alpha_{1}$ and $\alpha_{2}$ as $0.3580$ and $0.5454$, respectively, both significant at the $1\%$ level. The Wald test yields a chi-square statistic of $48.3438$ ($\text{p-value}<0.001$). In SHED, $\alpha_{1}$ and $\alpha_{2}$ are estimated as $0.3182$ and $0.6114$, respectively, both also significant at the 1% level, with a Wald test chi-square statistic of $111.5533$ ($\text{p-value}<0.001$). and with output-space distance [^16] In CLUE, our regression model estimates $\beta_{1}$ and $\beta_{2}$ as $0.0070$ and $0.0403$, respectively, both significant at the 1% level, with a Wald test chi-square statistic of $39.4285$ ($\text{p-value}<0.001$). In SHED, $\beta_{1}$ and $\beta_{2}$ are estimated as 0.0132-0.0132 and $0.0565$, respectively, again both significant at the 1% level, with a Wald test chi-square statistic of $116.2329$ ($\text{p-value}<0.001$)..
Regarding RQ3, while the CLUE vs. SHED condition itself does not have a direct impact on overall input distance ($\text{p-value}=0.442$), it significantly moderates the impact of aggregate distance and trajectory distance.  [^17] The interaction terms $\alpha_{4}=0.0399$ ($\text{p-value}=0.091$) and $\alpha_{5}=-0.0661$ ($\text{p-value}<0.001$) show that the effect of aggregate distance (and trajectory distance) on overall input-space distance becomes stronger (and weaker, respectively) in the CLUE (criminal history) condition than in the SHED (income) condition. A similar finding applies to output-space distance, considering $\beta_{4}=0.0201$ ($\text{p-value}<0.001$) and $\beta_{5}=-0.0163$ ($\text{p-value}<0.001$).
@@
After each input-space distance question, we have an optional question where participants select factor(s) that influence their assigned input-space distance: trajectory, aggregate, and “Other (please specify)”. Participants select trajectory more often than aggregate in both CLUE ($80.6\%>71.4\%$) and SHED ($75.2\%>60.0\%$) conditions, reaffirming our quantitative result that trajectory distance has a stronger impact than aggregate distance on overall input-space distance. We perform thematic coding inductively on the 55 free-text responses (rationales) that follow the third choice, “Other (please specify below)” ($1.9\%$ in CLUE, $4.6\%$ in SHED), and come up with $9$ themes (after filtering out $4$ vague rationales). Although we expect people to specify “Other” factors, most rationales are still about: (1) Both trajectory and aggregate ($23$ rationales); (2) Only trajectory ($12$ rationales); (3) Only aggregate ($3$ rationales).
New themes that influence how people evaluate overall input-space distance include: (4) Concern about unobserved factors ($4$ rationales, e.g., “Possible retirement or loss of job by B”); (5) First time step ($3$ rationales, e.g., “Similar/near equal high amount of arrest to start with.”); (6) Last time step ($2$ rationales, e.g., “In addition to the vastly different trajectories, it is important to consider the endpoint of the graph, as it is at this time that the decision will be made, such as by a bank.”); (7) Magnitude of acceleration ($2$ rationales, e.g., “Jumps in increased or decreased crimes.”). Two other rationales are unique enough to each define its own theme: (8) Comparing to an average person: “Although the total arrest times have a significant different, the number of total arrests for both people is quite a bit and more than an average person”; (9) Intersection: “The intersection of their trajectories”.
After each output-space distance question, participants may optionally explain their rationales, e.g., “How did you decide on this ‘fair’ risk score for A (compared to K)?”. We manually code rationales for the first set (out of three defendant/household sets) to minimize the impacts of survey fatigue (Porter, Whitcomb, and Weitzer 2004). After excluding vague, data-inconsistent, or AI-generated rationales, we have $243$ rationales (CLUE) and $194$ rationales (SHED).
Among those rationales, our “Only trajectory” code appears most often, assigned to $42.8\%$ of rationales in CLUE and $41.8\%$ in SHED (e.g., “They both are trending upward which makes me think they are similar. I gave them both the risk of $28\%$ regardless of the number of arrests.”). In contrast, our “Only aggregate” code is assigned to merely $21.0\%$ of rationales in CLUE and $15.5\%$ in SHED (e.g., “The graph shows that Household D has a significantly higher total income compared to Household K. Therefore, a fair risk score for Household D should be lower than $15\%$. […]”). This qualitative comparison reaffirms our quantitative finding that trajectory distance influences output-space distance more strongly than aggregate distance does, with the caveat that many rationales still factor in both trajectory and aggregate distances ($21.4\%$ in CLUE and $30.4\%$ in SHED), e.g., “Defendant D has been arrested a lot more than Defendant K, but he is trending down. That being said, I think it balances out and their likelihood of reoffending is probably similar.”
@@
One inertia metric that is relevant to both criminal justice and personal finance is childhood poverty, as explained earlier. However, for adults in our two real-world datasets, information on adults’ childhood socio-economic background is not available, especially not at the individual level. Therefore, to quantify inertia empirically, we may use race as an individual-level but imperfect proxy for childhood poverty (or inertia) in the U.S. as the National Center for Education Statistics (NCES) found that childhood poverty rate differs substantially across races (e.g., White: $13\%$, Asian: $14\%$, Pacific Islander: $25\%$, American Indian: $36\%$, Black: $39\%$) in 2012.181818https://nces.ed.gov/programs/coe/indicator/cce/family-characteristics—figure 4. Accessed Jan 19, 2025 To obtain $m$ per race, we scale these poverty rates by the maximum rate ($39\%$) so that $m$ spreads across $[0,1]$.
Given four equidistant time steps $\{0,1,2,3\}$ (assuming $\Delta t=1$), we model cumulative arrests in CLUE (or cumulative income, in multiples of \$10k\\$10k in SHED) at each time step as $X_{0}$, $X_{1}$, $X_{2}$, $X_{3}$. Their velocities are $V_{0}$, $V_{1}$, $V_{2}$, where $V_{i}=\frac{X_{i+1}-X_{i}}{\Delta t}=X_{i+1}-X_{i}$. Their accelerations are $A_{0}$ and $A_{1}$, where $A_{i}=\frac{V_{i+1}-V_{i}}{\Delta t}=V_{i+1}-V_{i}$. We average these two acceleration terms to get $A_{\text{avg}}=\frac{1}{2}(A_{0}+A_{1})=\frac{1}{2}(v_{1}+v_{2})$. Noting that the Effort-related feature is desirable in SHED (income) but undesirable in CLUE (arrests), we compute the individual-level (good) Effort, $E$, with context adaptations:
@@
The function $\sigma(z)=\frac{1}{1+e^{-z}}$ monotonically maps real values from $(-\infty,\infty)$ to the $(0,1)$ range, serving two purposes. First, it allows $\sigma(A_{\text{avg}})$ to obtain only one possible sign. More specifically, considering the desirable Effort feature, income, in SHED. Note that $m$ is always positive. When the average acceleration term is positive, $E=m\cdot A_{\text{avg}}$ will assign better (more positive) Effort to those with higher inertia (from more disadvantaged backgrounds). However, when the average acceleration term is negative, $E=m\cdot A_{\text{avg}}$ will assign worse (more negative) Effort to those with higher inertia, which contradicts our intention (given the same acceleration, more disadvantaged people should be credited for better Efforts). The sigmoid function eliminates the possibility of a negative sign and solves this problem. Second, the sigmoid function gathers extreme values closer together so that we have more data points to compute group parity even for subjects with extreme accelerations.
@@
To ensure the aggregate income feature stays in [0,1] like the Effort feature and reflect the intuition that the same income gap at lower income (e.g., \$10k\\$10k vs. \$20k\\$20k) is more significant than at higher income (e.g., \$110k\\$110k vs. \$120k\\$120k), we define an aggregate (sum) function $S(\mathbf{x})$ by applying a modified (scaled and translated) right-hand side of the sigmoid function $\sigma$ on the whole four-year (non-negative) income $X_{3}=\sum_{t=0}^{3}x_{t}$, scaled by $\lambda=\$200k$ so that this aggregate feature is empirically spread out across [0,1] range:
@@
Table 1 shows that according to our EaIF metric, XGBoost is less fair than the other two models, and random forest is slightly more fair than logistic regression. This result is consistent across both sets of weights (between the Effort feature and the most recent cumulative feature).
@@
Binns (2020) found that group fairness and individual fairness are rooted in similar philosophical concepts (“consistency” and “egalitarianism”). Algorithms that satisfy both group fairness and individual fairness were developed, demonstrating the two metrics’ compatibility (Zemel et al. 2013). Therefore, we will also demonstrate how to compute Effort-aware Group Fairness (EaGF) with our same force-based notion of “Effort”, which was validated by our human subjects experiment in the individual fairness context.
Given an individual-level Effort feature, our goal is to compute Effort-aware group parity and observe this metric as a function of numeric Effort. Our main steps include (1) partitioning decision subjects into similar-effort bins; (2) computing a demographic group parity metric within each bin; (3) plotting this parity metric against the Effort bins. We used demographic features of each household representative who completed the SHED survey to compute group fairness metrics.